An airplane travels 600 miles against the wind in 4 hours, and makes the return trip with the same wind in 2 hours. Find the speed of the wind.
Answer by TimothyLamb(4379)

x = wind speed
y = airplane speed in still air
---
s = d/t
---
against the wind:
y - x = 600/4
---
with the wind:
y + x = 600/2
---
put the system of linear equations into standard form
---
x - y = -150
x + y = 300
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = wind speed = 75 mph
y = airplane speed in still air = 225 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---